[The following information applies to the questions displayed below.] Glasgow Enterprises started the period with 70 units in beginning inventory that cost $2.70 each. During the period, the company purchased inventory items as follows: Purchase No. of Items Cost

1 220 $ 3.20

2 160 $ 3.30

60 $ 3.70

Glasgow sold 240 units after purchase 3 for $10.20 each. What is Glasgow's ending inventory under LIFO?
Answer #1

Total units available for sale= 70+220+160+60 = 510 units

Ending inventory = 510-240 = 270 units

Ending inventory under LIFO = 70*2.70 + 200*3.20

= $829